# Digital Twin-Empowered Deep Reinforcement Learning for Intelligent VNF Migration in Edge-Core Networks

### v1

@@ Abstract
The growing demand for services and the rapid deployment of virtualized network functions (VNFs) pose significant challenges for achieving low-latency and energy-efficient orchestration in modern edge–core network infrastructures. To address these challenges, this study proposes a Digital Twin (DT)-empowered Deep Reinforcement Learning framework for intelligent VNF migration that jointly minimizes average end-to-end (E2E) delay and energy consumption. By formulating the VNF migration problem as a Markov Decision Process and utilizing the Advantage Actor-Critic model, the proposed framework enables adaptive and real-time migration decisions. A key innovation of the proposed framework is the integration of a DT module composed of a multi-task Variational Autoencoder and a multi-task Long Short-Term Memory network. This combination collectively simulates environment dynamics and generates high-quality synthetic experiences, significantly enhancing training efficiency and accelerating policy convergence. Simulation results demonstrate substantial performance gains, such as significant reductions in both average E2E delay and energy consumption, thereby establishing new benchmarks for intelligent VNF migration in edge–core networks.
@@ I Introduction
Although DRL has demonstrated notable performance improvements, it is typically implemented to solve large action spaces in tasks such as VNF migration and requires a substantial amount of training data. This, in turn, necessitates frequent interactions with the physical environment to collect sufficient data, imposing considerable overhead on real-world systems. To mitigate the burden a Digital Twin (DT) can be implemented to simulate environment dynamics and generate additional traning data [11]. By serving as a virtual replica of the physical system, the DT enables the creation of a virtual experience buffer that, when combined with the physical expereince buffer enriches the training dataset and accelerates policy convergence while reducing real-world exploration costs.
@@
  • To enhance the performance of the DRL model, specifically the Advantage Actor-Critic (A2C) for intelligent VNF migration within edge–core networks, a novel integration with a DT is proposed. This integration significantly reduces the reliance on physical environment interactions, thereby lowering system overhead.
  • The DT module is composed of a multi-task Long Short-Term Memory (LSTM) network and a multi-task Variational Autoencoder (VAE). The multi-task LSTM continuously adapts to the environment by predicting state transition dynamics and reward feedback, while the multi-task VAE generates simulated state–action pairs. Together, the multi-task LSTM and VAE augment the training dataset by storing synthetic experiences in the DT experience buffer, thereby enhancing the training efficiency of the A2C model.
  • The proposed framework is applied to solve the joint optimization problem of minimizing both the E2E delay of VNF-FGs and the energy consumption in edge–core networks.
  • Extensive simulations demonstrate that the proposed framework significantly outperforms baseline methods. For example, in terms of average E2E delay and average energy consumption, the proposed framework achieves reductions of 7% and 15%, respectively, compared to baselines such as threshold-based VNF migration.
@@ II-B E2E Delay Model
where, $\tau$ denotes the baseline processing delay per data packet, and $\Gamma^{s}_{\text{CPU}}(t)$ is the CPU utilization of server $s$ at time step $t$, defined as:
@@ II-D Problem Formulation
The reduction in E2E delay resulting from VNF migration is denoted as:
$\delta(t)=\Upsilon-\Upsilon^{\prime}$ (14)
@@
Similarly, the reduction in energy consumption due to VNF migration is given by:
$\eta(t)=\Lambda-\Lambda^{\prime}$ (15)
@@
Optimization 1. For a set of $R$ VNF-FGs over time steps $t=1$ to $T$, the VNF migration problem in the edge–core network is formulated as the maximization of the total average reduction in E2E delay and energy consumption:
@@ III-B Description of the DRL Module
$\mathcal{R}(\mathcal{S}(t),\mathcal{A}(t))=sigm(\lambda),$ (22)
here, ${sigm}(\cdot)$ denotes the sigmoid function which is applied to normalize the reward within a bounded range.
@@ III-C Description of the DT Module
In this study, the training efficiency of the system is improved by establishing a DT of the physical environment. The digital twin focuses exclusively on modeling the environment’s feedback to the system’s actions, thereby simplifying its design and reducing the overhead typically associated with constructing a complete replica of the physical system. Within the DT framework, a multi-task VAE [15] is implemented to generate synthetic state–action pairs, while a multi-task LSTM [16] network is utilized to predict the corresponding next state and reward based on these outputs. This combination enables the DT to simulate realistic environment dynamics and augment the training process with high-quality synthetic data.
At time step $t$, the multi-task VAE model $\mathcal{V}$ takes as input the physical environment’s state $\mathcal{S}(t)$ and action $\mathcal{A}(t)$, and generates the corresponding state and action in the DT, denoted as $\mathcal{S}^{{DT}}(t)$ and $\mathcal{A}^{{DT}}(t)$. This process can be formally expressed as:
@@
The loss function for the multi-task VAE is composed of a reconstruction term and a regularization term. It can be expressed as:
@@
Moreover, at time step $t$, the multi-task LSTM model $\mathcal{T}$ utilizes the generated DT state $\mathcal{S}^{{DT}}(t)$ and action $\mathcal{A}^{{DT}}(t)$ as inputs to predict the next state $\mathcal{S}^{{DT}}(t+1)$ and the corresponding reward $\mathcal{R}^{{DT}}(t)$, and can be expressed as:
@@
The adopted loss function for the network is the mean squared error (MSE), can be formulated as:
@@
A DT experience buffer, denoted as $\mathcal{B}^{DT}$ is established to store the generated data tuples, comprising $\langle\mathcal{S}^{{DT}}(t),\mathcal{A}^{{DT}}(t),\mathcal{R}(\mathcal{S}^{{DT}}(t),\mathcal{A}^{{DT}}(t)),\mathcal{S}^{{DT}}(t+1)\rangle$. This virtual experience buffer is utilized alongside the physical experience buffers to train the A2C model, thereby improving sample efficiency and accelerating policy convergence.
@@ III-D DT-Empowered DRL Framework Training
To train the A2C model, the DT experience buffer $\mathcal{B}^{DT}$ and two physical experience buffers are utilized, namely, the successful physical experience buffer $\mathcal{B}^{P}_{{SUC}}$ and the unsuccessful physical experience buffer $\mathcal{B}^{P}_{{FAIL}}$. Separating successful and unsuccessful experiences helps the system distinguish between constraint-satisfying and constraint-violating actions, enabling more informed policy updates than conventional random sampling. At each time step $t$, after executing an action, if the constraints C1 and C2 are not satisfied, the corresponding experience is stored in $\mathcal{B}^{P}_{{FAIL}}$; otherwise, it is stored in $\mathcal{B}^{P}_{{SUC}}$. During each training iteration, the system randomly samples $\varkappa\cdot J^{P}$ experiences from $\mathcal{B}^{P}_{{SUC}}$ and $(1-\varkappa)\cdot J^{P}$ experiences from $\mathcal{B}^{P}_{{FAIL}}$ to form the physical mini-batch $M^{P}$. The parameter $\varkappa\in[0,1]$ controls the balance between successful and unsuccessful experiences and is critical for stable training. In the case of $\mathcal{B}^{{DT}}$, a DT mini-batch $M^{{DT}}$ is formed by randomly sampling $J^{{DT}}$ experiences from the $\mathcal{B}^{{DT}}$.
During the training iterations, the combined mini-batches $M^{P}$ and $M^{{DT}}$ are used to update the A2C model. It is important to note that the multi-task VAE and multi-task LSTM models within the DT are trained separately using experiences from $\mathcal{B}^{P}_{{FAIL}}$ and $\mathcal{B}^{P}_{{SUC}}$. The overall training and interaction process is illustrated in Figure 2.
@@ IV-A Simulation Setup
A physical edge–core network is generated based on the Waxman topology model with parameters $\alpha=0.5$ and $\beta=0.2$, following the configuration described in [17]. The network consists of 60 servers, with 20 designated as edge cloud servers and 40 as core cloud servers, as specified in [4]. The edge cloud servers are configured with 40 units of CPU resources and 16 units of memory resources, whereas the core cloud servers are configured with 200 units of CPU resources and 64 units of memory resources [4]. The CPU and memory utilization thresholds are both set to 80%. Additionally, the available bandwidth for each link is set to 3.5 Gbps [12]. The CPU and memory resource demands for each VNF within the VNF-FGs are randomly assigned, ranging from 1 to 20 units for CPU and from 1 to 4 units for memory, consistent with [4]. Regarding server energy consumption, the baseline energy consumption is set to 10 units, the maximum energy consumption is set to 110 units, and the transition energy consumption is set to 2 units. Each VNF-FG is composed of four VNFs. In each episode, 300 VNF-FG requests arrive sequentially, following a Poisson process with an average arrival rate of 0.2 requests per time step. The packet rate within each VNF-FG is set to 100 packets per second.
In the A2C model, both the actor and critic networks are implemented as fully connected neural networks with three hidden layers comprising 256, 128, and 64 neurons, respectively, all using ReLU activation. The actor network outputs deterministic actions through a Tanh-activated layer with 64 neurons, while the critic network outputs a single Q-value via a linear neuron. In the DT module, the multi-task LSTM network comprises a single LSTM layer with 64 hidden units, followed by a fully connected layer with 64 ReLU-activated neurons. Its output is split into two task-specific heads: one predicts the next state, the other the immediate reward, both using linear activation. The multi-task VAE module includes an encoder with two ReLU-activated hidden layers of 64 and 32 neurons, followed by two output layers estimating the mean and log-variance of the latent distribution. The decoder consists of a shared hidden layer with 64 ReLU-activated neurons and two task-specific Sigmoid heads for reconstructing the digital twin state and action, respectively. Table I summarizes the simulation parameters used for the A2C model and associated neural networks.
@@
Balance parameter ($\iota$) 0.35
@@
For benchmarking performance comparisons, two VNF migration methods are examined, (1) Threshold-based automatic VNF migration[9]: The system monitors the CPU and memory utilization on servers and triggers VNF migration when either resource exceeds predefined thresholds, migrating VNFs from overutilized servers to underutilized servers criteria such as current running load. (2) Random VNF migration: The system randomly migrates VNFs within the edge–core network without considering any optimization criteria.
@@ IV-B Numerical Results and Comparisons
Figure 3: (a) Avergae energy consumption vs. number of VNF-FGs, (b) Average E2E delay vs. number of VNF-FGs, (c) Normalized cumulative rewards vs. episodes.
Figure 3(a) illustrates that the proposed framework outperforms baseline approaches, such as the threshold-based and random scheme, in terms of average energy consumption across varying numbers of VNF-FGs. Specifically, it achieves reductions of 15% and 29%, respectively. This improvement is attributed to the integration of A2C with a DT, which enables effective management of a high-dimensional state–action space and leads to higher cumulative rewards and lower energy use. In contrast, both baseline methods activate a larger number of servers during VNF migration, resulting in increased energy consumption.
Figure 3(b) presents the average E2E delay of the proposed framework compared to baseline approaches, such as the threshold-based and random scheme. The proposed method achieves the lowest average E2E delay, with reductions of 7% and 18%, respectively. This improvement stems from the reward function design in A2C, which jointly optimizes reductions in energy consumption and E2E delay before and after VNF migration.
Figure 3(c) shows that the normalized cumulative reward of the A2C model in the proposed framework achieves significantly faster convergence, stabilizing in as few as 12 episodes. In contrast, the baseline approaches, such as the threshold-based and random scheme exhibit no convergence behavior, as they are not learning-based methods.
@@ V Conclusions
In this study, a novel Digital Twin (DT)-assisted Deep Reinforcement Learning (DRL) framework is proposed to enhance the efficiency and intelligence of VNF migration in edge–core networks. By integrating the Advantage Actor-Critic model with a DT that captures network dynamics through a multi-task Long Short-Term Memory network and a multi-task Variational Autoencoder, the proposed solution effectively achieves the objective of reducing both average end-to-end delay and energy consumption. The incorporation of synthetic experience buffers alongside physical experience buffers further improves sample efficiency, enabling robust training even under highly dynamic workload conditions. Extensive evaluation demonstrates that the proposed framework consistently outperforms traditional threshold-based and random migration strategies across key performance metrics, such as average energy consumption, average E2E delay, and convergence speed. These findings validate the effectiveness of leveraging DT and DRL to enable scalable, adaptive, and sustainable orchestration of VNFs.

### v2

@@ Abstract
The growing demand for services and the rapid deployment of virtualized network functions (VNFs) pose significant challenges for achieving low-latency and energy-efficient orchestration in modern edge–core network infrastructures. To address these challenges, this study proposes a Digital Twin (DT)-empowered Deep Reinforcement Learning framework for intelligent VNF migration that jointly minimizes average end-to-end (E2E) delay and energy consumption. The VNF migration task is formulated as a Markov Decision Process and optimized using an Advantage Actor–Critic architecture, enabling adaptive migration decisions under dynamic network conditions. A key innovation of the proposed framework is the integration of a DT module composed of a multi-task variational autoencoder serving as the generative model and a multi-task long short-term memory network serving as the predictive model. Together, these models simulate environment dynamics and generate high-quality synthetic experiences that enhance training efficiency and accelerate policy convergence. Simulation results demonstrate substantial improvements in both end-to-end delay and energy consumption, establishing a new benchmark for intelligent VNF migration in edge–core networks.
@@ I Introduction
Although DRL has demonstrated notable performance improvements, it is typically implemented to solve large action spaces in tasks such as VNF migration and requires a substantial amount of training data. This, in turn, necessitates frequent interactions with the physical environment to collect sufficient data, imposing considerable overhead on real-world systems. To mitigate the burden a Digital Twin (DT) can be implemented to simulate environment dynamics and generate additional training data [15]. By serving as a virtual replica of the physical system, the DT enables the creation of a virtual experience buffer that, when combined with the physical experience buffer enriches the training dataset and accelerates policy convergence while reducing real-world exploration costs. Furthermore, hosting DRL training processes within the DT can offload computational workloads from the physical system, thereby reducing energy consumption and minimizing disruptions caused by continuous model training.
@@
  • To minimize the E2E delay of VNF-FGs and reduce the overall energy consumption of edge–core networks, this work proposes a DRL framework based on an Advantage Actor–Critic (A2C) architecture that intelligently migrates VNFs within the edge–core infrastructure.
  • To enhance the performance and sustainability of the DRL framework for intelligent VNF migration, a novel integration with a DT is proposed that generates synthetic experience data and hosts the A2C training process. By shifting training from the physical environment to the DT, the framework significantly reduces reliance on real-world interactions and minimizes the energy cost and operational disruptions associated with continuous model updates.
  • The DT module incorporates a multi-task Variational Autoencoder (VAE) as the generative model and a multi-task Long Short-Term Memory (LSTM) network as the predictive model. The generative model first produces simulated state–action pairs, and the predictive model then uses these to generate the corresponding simulated next states and reward values. Together, these components enrich the training dataset with synthetic transitions stored in the DT experience buffer, thereby improving the training efficiency of the A2C model.
  • Extensive simulations demonstrate that the proposed framework significantly outperforms baseline methods. For example, in terms of average E2E delay and average energy consumption, the proposed framework achieves reductions of 7% and 15%, respectively, compared to baselines such as non DT-based A2C VNF migration.
@@ II-B E2E Delay Model
where, $\tau$ denotes the baseline processing delay per flow, and $\Gamma^{s}_{\text{CPU}}(t)$ is the CPU utilization of server $s$ at time step $t$, defined as:
@@ II-D Problem Formulation
The normalized reduction in E2E delay resulting from VNF migration is denoted as:
$\delta(t)=\frac{\Upsilon-\Upsilon^{\prime}}{\Upsilon}$ (14)
@@
Similarly, the normalized reduction in energy consumption due to VNF migration is given by:
$\eta(t)=\frac{\Lambda-\Lambda^{\prime}}{\Lambda}$ (15)
@@
Optimization 1. For a set of $R$ VNF-FGs over time steps $t=1$ to $T$, the VNF migration problem in the edge–core network is formulated as the maximization of the total average normalized reduction in E2E delay and energy consumption:
@@ III-B Description of the DRL Framework
$\mathcal{R}(\mathcal{S}(t),\mathcal{A}(t))=tanh(\lambda),$ (22)
here, ${tanh}(\cdot)$ denotes the hyperbolic tangent function which is applied to normalize the reward within a bounded range.
@@ III-C Description of the DT Module
In this study, a DT is established to enhance the performance and sustainability of the A2C model. Within the DT module, a multi-task VAE is implemented as the generative model to produce synthetic state–action pairs, and a multi-task LSTM network is utilized as the predictive model to estimate the corresponding next states and rewards. This combination enables the DT to emulate realistic environment dynamics and augment the learning process with high-quality synthetic data. Furthermore, by shifting the A2C training process from the physical environment to the DT, the framework significantly lowers the energy overhead and mitigates operational disruptions caused by frequent model updates.
At time step $t$, the generative model $\mathcal{V}$ takes as input the physical environment’s state $\mathcal{S}(t)$ and action $\mathcal{A}(t)$, and generates the corresponding state and action in the DT, denoted as $\mathcal{S}^{{DT}}(t)$ and $\mathcal{A}^{{DT}}(t)$. This process can be formally expressed as:
@@
The loss function for the generative model is composed of a reconstruction term and a regularization term. It can be expressed as:
@@
Moreover, at time step $t$, the predictive model $\mathcal{T}$ utilizes the generated DT state $\mathcal{S}^{{DT}}(t)$ and action $\mathcal{A}^{{DT}}(t)$ as inputs to predict the next state $\mathcal{S}^{{DT}}(t+1)$ and the corresponding reward $\mathcal{R}^{{DT}}(t)$, and can be expressed as:
@@
The adopted loss function for the predictive model is the mean squared error (MSE), can be formulated as:
@@
A DT experience buffer, denoted as $\mathcal{B}^{DT}$ is established to store the generated data tuples, comprising $\langle\mathcal{S}^{{DT}}(t),\mathcal{A}^{{DT}}(t),\mathcal{R}(\mathcal{S}^{{DT}}(t),\mathcal{A}^{{DT}}(t)),\mathcal{S}^{{DT}}(t+1)\rangle$. This virtual experience buffer is utilized alongside the physical experience buffers to train the A2C, thereby improving sample efficiency and accelerating policy convergence.
@@ III-D DT-Empowered DRL Framework Training
To train the A2C, the DT experience buffer $\mathcal{B}^{DT}$ and two physical experience buffers are utilized, namely, the successful physical experience buffer $\mathcal{B}^{P}_{{SUC}}$ and the unsuccessful physical experience buffer $\mathcal{B}^{P}_{{FAIL}}$. Separating successful and unsuccessful experiences helps the system distinguish between constraint-satisfying and constraint-violating actions, enabling more informed policy updates than conventional random sampling. At each time step $t$, after executing an action, if the constraints C1 and C2 are not satisfied, the corresponding experience is stored in $\mathcal{B}^{P}_{{FAIL}}$; otherwise, it is stored in $\mathcal{B}^{P}_{{SUC}}$. During each training iteration, the system randomly samples $\varkappa\cdot J^{P}$ experiences from $\mathcal{B}^{P}_{{SUC}}$ and $(1-\varkappa)\cdot J^{P}$ experiences from $\mathcal{B}^{P}_{{FAIL}}$ to form the physical mini-batch $M^{P}$. The parameter $\varkappa\in[0,1]$ controls the balance between successful and unsuccessful experiences and is critical for stable training. In the case of $\mathcal{B}^{{DT}}$, a DT mini-batch $M^{{DT}}$ is formed by randomly sampling $J^{{DT}}$ experiences from the $\mathcal{B}^{{DT}}$.
During the training iterations, the combined mini-batches $M^{P}$ and $M^{{DT}}$ are used to update the A2C actor and critic networks. After training, the updated actor parameters are transferred back to the physical environment for action selection. It is important to note that the generative and predictive models within the DT are trained using experiences drawn from $\mathcal{B}^{P}_{{FAIL}}$ and $\mathcal{B}^{P}_{{SUC}}$ at the beginning of each episode. The overall training and interaction process is illustrated in Figure 2.
@@ IV-A Simulation Setup
A physical edge–core network is generated based on the Waxman topology model with parameters $\alpha=0.5$ and $\beta=0.2$, following the configuration described in [11]. The network consists of 100 servers, with 35 designated as edge cloud servers and 65 as core cloud servers, as specified in [14]. The edge cloud servers are configured with 40 units of CPU resources and 16 units of memory resources, whereas the core cloud servers are configured with 200 units of CPU resources and 64 units of memory resources [14]. The CPU and memory utilization thresholds are both set to 80%. Additionally, the available bandwidth for each link is set to 3.5 Gbps [13]. The CPU and memory resource demands for each VNF within the VNF-FGs are randomly assigned, ranging from 1 to 20 units for CPU and from 1 to 4 units for memory, consistent with [14]. Each VNF-FG is composed of four VNFs. In each episode, 300 VNF-FG requests arrive sequentially, following a Poisson process with an arrival rate of 0.2 requests per time step.
In the A2C model, both the actor and critic networks are implemented as fully connected neural networks with four hidden layers comprising 1024, 512, 256, and 128 neurons, respectively, each using ReLU activation. The actor network generates a probability distribution over all candidate migration actions through a Softmax output layer. The critic network estimates the state-value using a single linear output neuron. In the DT module, the predictive model comprises a single LSTM layer with 64 hidden units, followed by a fully connected layer with 64 ReLU-activated neurons. Its output is split into two task-specific heads: one predicts the next state, the other the immediate reward, both using linear activation. The generative model includes an encoder with two ReLU-activated hidden layers of 64 and 32 neurons, followed by two output layers estimating the mean and log-variance of the latent distribution. The decoder consists of a shared hidden layer with 64 ReLU-activated neurons and two task-specific Sigmoid heads for reconstructing the digital twin state and action, respectively. Table I summarizes the simulation parameters used for the A2C and associated neural networks.
@@
Balance parameter ($\varkappa$) 0.30
@@
For benchmarking performance comparisons, two VNF migration schemes are evaluated: (1) a non–DT-based A2C VNF migration and (2) a random VNF migration.
@@ IV-B Numerical Results and Comparisons
Figure 3: (a) Average energy consumption vs. number of VNF-FGs, (b) Average E2E delay vs. number of VNF-FGs, (c) Average request acceptance ratio vs. number of VNF-FGs.
Figure 3(a) illustrates that the proposed framework outperforms baseline approaches, such as the non–DT-based A2C VNF migration and random VNF migration scheme, in terms of average energy consumption across varying numbers of VNF-FGs. Specifically, it achieves reductions of 15% and 27%, respectively. This gain results from integrating A2C with the DT module, which provides synthetic constraint-aware transitions that improve the A2C’s ability to identify energy-efficient migration decisions. By exposing the policy to a broader range of system states, the DT helps avoid unnecessary server activations and accelerates convergence toward energy-minimizing behavior.
Figure 3(b) presents the average E2E delay of the proposed framework compared to baseline approaches, such as the non–DT-based A2C VNF migration and random VNF migration scheme. The proposed method achieves the lowest average E2E delay, with reductions of 7% and 18%, respectively. This improvement stems from the DT-generated constraint-aware synthetic transitions, which offer richer exposure to delay-critical states. Consequently, the policy avoids delay inducing actions such as migrating VNFs to overloaded servers and converges toward more delay-efficient migration strategies.
Figure 3(c) illustrates that the proposed framework achieves a higher average request acceptance ratio compared to both the non–DT-based A2C VNF migration and the random VNF migration scheme. This improvement is attributed to the DT–assisted learning process which enables the policy to anticipate resource contention and avoid infeasible migration decisions by leveraging synthetic, constraint-aware experiences, thereby maintaining higher service admission rates under dynamic network conditions.
@@ V Conclusions
In this study, a novel DT-assisted DRL framework is proposed to enhance the efficiency of VNF migration in edge–core networks. By integrating the A2C model with a DT that captures network dynamics through a multi-task LSTM network and a multi-task VAE, the proposed solution effectively achieves the objective of minimizing both average end-to-end delay and energy consumption. The incorporation of synthetic experience buffers alongside physical experience buffers further improves sample efficiency, enabling robust training even under highly dynamic workload conditions. Extensive evaluation demonstrates that the proposed framework consistently outperforms traditional non–DT-based VNF migration and random VNF migration schemes across key performance metrics, such as average energy consumption, average E2E delay, and acceptance ratio. These findings validate the effectiveness of leveraging DT and DRL to enable scalable, adaptive, and sustainable orchestration of VNFs.